The species mass densities of a three?component (A, B, and C) liquid mixture are: acetone, \rho A= 326.4, kg/m3, acetic acid, \rho B = 326.4 kg/m3, and ethanol, \rho C = 217.6 kg/m3.

Determine the following for this mixture:
1. The mass fraction of each species in the mixture.

2. The mole fraction of each species in the mixture.

3. The mass of each component required to make one cubic meter of mixture.
Answer #1

Given:

The species mass densities of a three‐component (A, B, and C) liquid mixture are: acetone, \rhoA= 326.4, kg/m3, acetic acid, \rhoB = 326.4 kg/m3, and ethanol, \rhoC = 217.6 kg/m3.

Assume 1 m3 of total mixture

1. mass fraction of each mixture:

Mass of acetone in the mixture WA = 326.4 kg

Mass of acetic acid in the mixtureWB= 326.4 kg

Mass of ethanol in the mixtureWC= 217.6 kg

then, Total mass of the mixture will be WT = 870.4 kg

Mass frsction of acetone mA = WA/WT = 326.4/870.4 = 0.375

Mass fraction of acetic acid mB = WB/WT= 326.4/870.4= 0.375

Mass fraction of ethanol mC= WC/WT= 217.6/870.4= 0.25

2. Mole fraction of each component:

Molar mass of acetone MA = 58.08 g/mol

Molar mass of acetic acid MB = 60.05 g/mol

Molar mass of ethanol MC= 46.068 g/mol

Calculate of no. moles of each component:

Moles of acetone nA = WA/MA = 326.4/58.08 = 5.62 mol

Moles of acetic acid nB = 326.4/60.05= 5.44 mol

Moles of ethanol nC= WC/MC = 217.6/46.068= 4.72 mol

Total no of moles nT = 15.78 mol

mole fraction of acetone xA = nA/nT = 5.62/15.78= 0.356

mole fraction of acetic acid xB= nB/nT= 5.44/15.78= 0.345

mole fraction of ethanol xC= nC/nT= 4.72/15.778= 0.299

3. Mass required of each component to make 1 m3 of mixture will be=density of component * 1 m3

   Hence mass of each component required will be 326.4 kg of acetone, 326.4 kg of acetic acid & 217.6 kg of ethanol.